# MCPVerse: An Expansive, Real-World Benchmark for Agentic Tool Use

### v1

@@ Abstract
Large Language Models (LLMs) are evolving from text generators into reasoning agents. This transition makes their ability to use external tools a critical capability. However, evaluating this skill presents a significant challenge. Existing benchmarks are often limited by their reliance on synthetic tools and severely constrained action spaces. To address these limitations, we introduce MCPVerse, an expansive, real-world benchmark for evaluating agentic tool use. MCPVerse integrates more than 550 real-world, executable tools to create an unprecedented action space exceeding 140k tokens, and employs outcome-based evaluation with real-time ground truth for time-sensitive tasks. We benchmarked the state-of-the-art LLMs across three modes (Oracle, Standard, and Max-Scale), revealing that while most models suffer performance degradation when confronted with larger tool sets, the agentic models, such as Claude-4-Sonnet, can effectively leverage expanded exploration spaces to improve accuracy. This finding not only exposes the limitations of state-of-the-art models in complex, real-world scenarios but also establishes MCPVerse as a critical benchmark for measuring and advancing agentic tool use capabilities.
Codehttps://github.com/hailsham/mcpverse
Introduction
@@
First, many benchmarks rely on artificial tools (Patil et al. 2025; Li et al. 2023b). These often simulate calculators, simplified weather services, or mock shopping carts, whose data formats and interaction patterns bear little resemblance to production systems. This discrepancy allows models to succeed by recognizing superficial patterns rather than demonstrating the robust planning and coordination skills required for real-world tasks. While some benchmarks (Qin et al. 2024) claim to incorporate a wide range of real-world APIs, they often stop short of actual execution due to the complexities of live integration. Evaluation is thus confined to assessing the correctness of the selected tool name and its parameters, rather than the functional outcome of the interaction.
Second, existing benchmarks severely constrain the action space available to the model during evaluation. Even when a large set of APIs is listed, context length limitations force designers to mount only a small subset, often relying on a retrieval module to select a few dozen relevant options per query (Fei, Zheng, and Feng 2025; Li et al. 2023b; Qin et al. 2024). This strategy, while keeping prompts within the token limits of current models, prevents the assessment of a model’s ability to navigate a vast and complex solution space.
@@
  • Expansive Action Space: We have carefully curated a collection of $65$ MCPs, encompassing $552$ unique tools that cover a diverse range of functionalities: file system operations, version control (Git), financial data (Yahoo Finance), news aggregation (GeekNews), lifestyle services (Amap, Variflight), office productivity (Excel), and code sandbox. The combined schemas of these tools exceed $140,000$ tokens, surpassing the context and tool-mounting limits of many state-of-the-art models and providing an unprecedentedly large action space for exploration.
  • Hybird Outcome-Based Evaluation: Recognizing that a single user request can have multiple valid solution paths, our evaluation focuses on the final outcome rather than the specific sequence of tools used. We do not penalize models for deviating from a prescribed path. We employ a hybrid evaluation method: an LLM-as-a-judge assesses the correctness of text-based outputs, while automated evaluation scripts verify state changes for tasks involving file system modifications or other environmental interactions.
Figure 1: The MCPVerse Evaluation System. A user’s task is processed by an LLM agent that can access a diverse toolset built upon MCP pool collected from MCP hubs. The agent autonomously plans and executes a sequence of function calls, and its final response is checked for correctness by an automated evaluation system.
Using our evaluation suite, we benchmarked 8 leading LLMs. Our findings show that the top-performing model, Claude-4-Sonnet, achieved an accuracy of only $57.77\%$, indicating significant room for improvement across the field. Furthermore, our experiments revealed that many state-of-the-art models are not yet equipped to handle the scale of MCPVerse, facing limitations such as a 64K context length (DeepSeek-V3), a 128-tool limit (GPT-4o-20241120), or a 512-tool limit (Gemini-2.5-Pro).
To accommodate these constraints while still probing performance at scale, we designed three evaluation modes: Oracle mode (only the pre-annotated required MCPs are provided), Standard mode (all potentially relevant MCPs are provided), and Max-Scale Mode (all 65 MCPs are mounted). Unsurprisingly, most models exhibited performance degradation as the number of available MCPs increased. More strikingly, Claude-4-Sonnet achieved a higher score in Standard mode than in the simpler Oracle mode. This counter-intuitive result underscores the significant benefit of a larger exploration space, which allows more capable models to discover more effective or robust solutions.
@@
  • We introduce MCPVerse benchmark, a comprehensive benchmark built on a large-scale set of executable, real-world tools. The benchmark features meticulously designed tasks, each accompanied by annotated ground truth or a script for time-sensitive queries.
  • We build an end-to-end, automated evaluation system that facilitates multi-turn interactions between the LLM agent and MCP tools. The system assesses the final outcome using our hybrid, outcome-based metric, which combines an LLM-as-a-judge for textual answers with dedicated scripts for verifying environmental state changes.
@@
The benchmark and evaluation code will be made publicly available.
Related Work
Agentic Tool Use
There is a growing interest in equipping LLMs with external tools to overcome their inherent limitations in real-world applications. Fine-tuned models such as Toolformer (Schick et al. 2023), ToolLLaMA (Qin et al. 2024), and Gorilla (Patil et al. 2024) have demonstrated the ability to accurately use different tools like calculators and search engines. In parallel, strategies such as ToolkenGPT (Hao et al. 2023), METATOOL (Wang et al. 2024b), and IPR (Xiong et al. 2024) have been introduced to enhance the compatibility and effective use of various tools. Furthermore, agentic frameworks like WebMap (Spiegel and Horák 2024), ReAct (Yao et al. 2023), Voyager (Wang et al. 2024a), Agent Reasoning (Wu, Zhu, and Liu 2025), Middleware (Gu et al. 2024), and ViperGPT (Sur´ıs, Menon, and Vondrick 2023) enable models to perform complex, multi-step tasks. These applications range from online shopping and software repair to handling multimodal inputs and learning skills in interactive environments like Minecraft.
Tool-Use Benchmarks
@@
Numerous benchmarks have been developed to evaluate the tool-use capabilities of LLMs, each contributing different methodologies. Some efforts, such as ToolAlpaca (Tang et al. 2023) and NexusRaven (Srinivasan et al. 2023), have focused on automatic data generation to create large-scale evaluation sets. Another line of work has concentrated on creating diverse execution environments. For instance, Toolbench (Qin et al. 2024) provides a collection of non-executable APIs, while a more common strategy involves building sandboxed environments to simulate real-world interactions. BFCL-v3 (Patil et al. 2024), for example, implements codebases to simulate services like Twitter and mathematical calculators, and API-Bank (Li et al. 2023b) simulates common tasks such as sending emails or querying stock prices with mock data. Focusing on mobile-centric scenarios, ToolSandbox (Lu et al. 2025) simulates tools for managing device states like Wi-Fi, cellular service, and location service, while HammerBench (Wang et al. 2025) replicates API functionalities based on the features of 60 commercial applications. Other benchmarks have concentrated on specific domains, with ComplexFuncBench (Zhong et al. 2025) targeting travel-related tasks, $\tau$-bench (Yao et al. 2024) focusing on retail and airline scenarios, and ACEBench (Chen et al. 2025) covering four scenarios including food delivery and financial services. Fundamentally, these benchmarks assess model capabilities using simulated, simplified tools; their interaction logic and return values are based on fixed rules or templates, lacking the dynamism and uncertainty of real-world environments, and the complexity of the data they handle does not match that of practical applications.
@@
While previous work has attempted to connect LLMs with real-world APIs, such as in AutoFeedback (Liu et al. 2024; Song et al. 2023), these efforts often faced challenges due to complex and non-standardized interfaces. The Model Context Protocol (MCP) (Anthropic 2024) was introduced to address this issue. MCP is a standardized protocol for LLM-tool interaction that enables dynamic discovery and orchestration of tools based on task requirements. As the protocol has matured, it has seen widespread adoption across the industry(OpenAI 2025; cur 2025; Anthropic 2025). Consequently, recent work has begun to establish benchmarks on top of MCP. For instance, MCP-Zero (Fei, Zheng, and Feng 2025) proposed a retrieval strategy for MCPs and collected an associated toolset, MCP-Tools. Meanwhile, MCPBench (Luo et al. 2025) focused on comparing the performance of different MCPs for web-search tasks.
Finally, existing benchmarks are constrained by the number of tools they can support concurrently and their restriction to non-real-time tasks. Consequently, they fail to capture the dynamic, adaptive, and complex nature of real-world agentic interactions. To address these limitations, we introduce MCP Verse, a comprehensive benchmark designed to simulate realistic, multi-step agentic workflows. MCP-Verse provides a robust evaluation framework by incorporating outcome-based evaluation, full MCP compatibility, real-time tasks, and concurrent support for up to 552 tools, thereby bridging the gap between benchmarking and real-world agent performance.
MCPVerse Design
Figure 2: Example tasks from MCPVerse. The figure shows tasks at three complexity levels, each with annotations for the required tools and the ground truth answer, which can be either static or generated by a script for time-sensitive questions.
MCP Collection
@@
We collect a diverse set of MCPs from platforms such as Smithery, mcp.so, and ModelScope. These MCPs offer a wide range of functionalities, including local file operations, real-time flight booking, and SQL querying. As a Results, we carefully selected MCPs to be included in our benchmark based on the following criteria.
@@
In total, the selected $65$ MCPs expose more than $552$ individual tools. This number may fluctuate, as MCP providers have the discretion to add or deprecate tools over time. The configuration files for all MCPs used in our benchmark are included in our codebase.
Data Curation
@@
The collected MCP collection defines a large and diverse action space, enabling a wide variety of task types to be executed. Based on this action space, we construct a benchmark comprising $250$ tasks, which are categorized into three levels of complexity based on their required tool orchestration pattern. Figure 3 illustrates the distribution of these tasks by both complexity and time-sensitivity.
  • L1: These tasks can be completed by a single tool within $1$ or $2$ steps.
  • L2: These tasks necessitate a solution pathway of at least $5$ steps and may involve the use of a single tool or multiple tools.
  • L3: These tasks are relatively complex and may require the collaboration of different tools or the in-depth application of a specific tool, and usually take a minimum of more than 5 steps to solve.
Figure 3: Questions distribution
@@
During the task construction phase, annotators were instructed to select interested MCPs from MCP Hubs by following the standards specified in the previous section, and then construct tasks based on these MCPs. They were required to carefully understand the definitions of each tool within the MCPs. The constructed tasks needed to meet the following characteristics: (1) There must be a solution path for the task within the selected set of MCPs; (2) The answers should be as objective and unambiguous as possible to facilitate LLM scoring; (3) The tasks should preferably be close to real-world scenarios. (4) The task should be unsolvable by the model without the use of external tools. Annotators used some MCP Client tools (e.g., Trae) to assist in question generation and answer retrieval. Each task was required to be annotated with the following attributes: question, required MCPs, required tools, time-sensitivity, complexity, and ground truth. Following the initial construction, every task underwent a comprehensive review. During this stage, we authored custom scripts to retrieve real-time answers for all time-sensitive questions.
The Choice of No Retriever
Unlike many previous tool-use benchmarks that incorporate a retrieval module (Fei, Zheng, and Feng 2025; Li et al. 2023b; Qin et al. 2024), our evaluation system deliberately omits this component for two primary reasons.
First, we hypothesize that agentic models can discover novel solution paths when presented with a large-scale action space. A retriever trained on existing patterns might inadvertently hinder this crucial exploratory potential.
Second, we anticipate a paradigm shift where LLMs, powered by the growing adoption of MCP, will function as operating systems. In this future, the number of tools connected to an LLM will vastly exceed the dozens typically seen in current benchmarks. For comparison, the average smartphone has about $80$ installed apps (buildfire 2025) (by the way, many of these apps begin to offer their own MCP services), and considering each app exposes multiple APIs (analogous to tools), the total number of available tools could easily reach the hundreds or thousands. Even with a retriever, the scale of the candidate toolset an LLM must handle will be far greater than what is benchmarked today.
Metric
Given MCPVerse’s expansive action space, there can be multiple ways to fulfill the demand, which we are not able to measure, but the end result, or end state, should be consistent with ground truth labelling. Therefore we adopt a hybrid approach for this: for tasks with textual ground truth, an LLM-as-a-judge assesses the response’s correctness, while for other tasks, such as verifying file modifications, a dedicated evaluation script is used. The scoring model we use for the LLM-as-a-judge is GPT-4o-20241120. We note that using open-source models, such as Qwen-32B, also yields a consistent ranking of results (see Appendix 2 for more details). All scoring is binary, marking the final result as either correct or incorrect.
Evaluation Results
Setup
Our evaluation system (Figure 1) is built upon the CAMEL framework (Li et al. 2023a). We aggregates all available tools from MCPs and presents them to the LLMs through a standardized function-calling mechanism. However, this unified approach introduces two significant challenges: model context length limits and navigating API-specific restrictions.
The first and most significant challenge is context length. The full MCPVerse benchmark includes $552$ tools, and their combined schemas require approximately 147k tokens. This size exceeds the context capacity of many leading models, including DeepSeek-V3 (64k), GPT-4o-20241120 (128k), and Qwen3-235B-A22B (128k). Only models with exceptionally large context windows, such as Claude-4-Sonnet (200k) and Gemini-2.5-Pro (1M), can process the entire toolset at once.
To ensure a fair and comprehensive evaluation across models, we designed three distinct modes:
  • Oracle Mode: In this setting, we load only the minimal set of MCPs required to solve a given problem. While multiple solution paths might exist, we annotate one feasible path for each question and mount only the corresponding MCPs. The context length therefore varies for each question.

### v2

@@ Abstract
Large Language Models (LLMs) are evolving from text generators into reasoning agents. This transition makes their ability to use external tools a critical capability. However, evaluating this skill presents a significant challenge. Existing benchmarks are often limited by their reliance on synthetic tools and severely constrained action spaces. To address these limitations, we introduce MCPVerse, an expansive, real-world benchmark for evaluating agentic tool use. MCPVerse integrates more than 550 real-world, executable tools to create an unprecedented action space exceeding 147k tokens, and employs outcome-based evaluation with real-time ground truth for time-sensitive tasks. We benchmarked the state-of-the-art LLMs across three modes (Oracle, Standard, and Max-Scale), revealing that while most models suffer performance degradation when confronted with larger tool sets, the agentic models, such as Claude-4-Sonnet, can effectively leverage expanded tool spaces to improve accuracy. This finding not only exposes the limitations of state-of-the-art models in complex, real-world scenarios but also establishes MCPVerse as a critical benchmark for measuring and advancing agentic tool use capabilities. The code and dataset are avaiable at https://github.com/hailsham/mcpverse.
1 Introduction
@@
First, many benchmarks rely on artificial tools (Patil et al., 2025; Li et al., 2023b). These often simulate calculators, simplified weather services, or mock shopping carts, whose data formats and interaction patterns bear little resemblance to production systems. This discrepancy allows models to succeed by recognizing superficial patterns rather than demonstrating the robust planning and coordination skills required for real-world tasks. While some benchmarks (Qin et al., 2024) claim to incorporate a wide range of real-world APIs, they often stop short of actual execution due to the complexities of live integration. Evaluation is thus confined to assessing the correctness of the selected tool name and its parameters, rather than the functional outcome of the interaction.
Second, existing benchmarks severely constrain the action space available to the model during evaluation. Even when a large set of APIs is listed, context length limitations force designers to mount only a small subset, often relying on a retrieval module to select a few dozen relevant options per query (Fei et al., 2025; Li et al., 2023b; Qin et al., 2024). This strategy, while keeping prompts within the token limits of current models, prevents the assessment of a model’s ability to navigate a vast and complex solution space.
@@
  • Expansive Action Space: We have carefully curated a collection of $65$ MCPs, encompassing $552$ unique tools that cover a diverse range of functionalities. The combined schemas of these tools exceed $147k$ tokens, surpassing the context and tool-mounting limits of many state-of-the-art models and providing an unprecedentedly large tool space for exploration and exploitation.
  • Hybird Outcome-Based Evaluation: Recognizing that a single task can have multiple valid solution paths, our evaluation focuses on the final outcome rather than the specific sequence of tools used. We do not penalize models for deviating from a prescribed path. We employ a hybrid evaluation method: an LLM judges semantic consistency between the outcome and the ground truth, while automated evaluation scripts verify state changes for tasks involving file system modifications or other environmental interactions.
Using our evaluation system, we benchmarked 12 leading LLMs. Our findings show that the top-performing model, Claude-4-Sonnet, achieved an success rate of only $44.2$ at max-scale mode, indicating significant room for improvement across the field. Furthermore, our experiments revealed that many state-of-the-art models are not yet equipped to handle the scale of MCPVerse, facing limitations such as a 64K context length (DeepSeek-V3), a 128-tool limit (GPT-4o-20241120), or a 512-tool limit (Gemini-2.5-Pro).
To accommodate these constraints while still probing performance at scale, we designed three evaluation modes: Oracle mode (only per-question minimal MCPs are provided), Standard mode (all potentially relevant MCPs are provided), and Max-Scale Mode (all 65 MCPs are mounted). Unsurprisingly, most models exhibited performance degradation as the number of available MCPs increased. More strikingly, agentic models like Claude-4-Sonnet, Qwen3-235B-2507 and GLM-4.5 achieved a higher score in Standard mode than in the simpler Oracle mode. This counter-intuitive result suggests that a larger tool space enables emergent solution paths and “hacking” behaviors, allowing capable models to uncover alternatives that are inaccessible under constrained settings.
@@
  • We introduce MCPVerse benchmark, a comprehensive benchmark built on a large-scale set of executable, real-world tools. The benchmark features meticulously designed tasks, each annotated with ground truth (or an evaluation script for time-sensitive queries), as well as additional metadata such as required MCPs, complexity level, and task type.
  • We build an end-to-end, automated evaluation system that facilitates multi-step interactions between the LLM agent and MCP tools. The system assesses the final outcome using our hybrid, outcome-based metric, which combines an LLM-as-a-judge for textual answers with dedicated scripts for verifying environmental state changes.
@@
2 Related Work
Figure 1: The Overview of MCPVerse. MCPVerse evaluation system. Tasks are curated with metadata and associated data files of various formats. Tools are collected from MCP hubs into an MCP pool and mounted as a toolset under three modes: Oracle, Standard, and Max-Scale. An LLM agent executes an agentic loop—planning and invoking tools via function calls—and produces an outcome. For time-invariant tasks the outcome is compared with human annotations; for time-sensitive tasks real-time scripts retrieve ground truth; an automated scorer then determines correctness.
2.1 Agentic Tool Use
There is a growing interest in equipping LLMs with external tools to overcome their inherent limitations in real-world applications. Fine-tuned models such as Toolformer (Schick et al., 2023), ToolLLaMA (Qin et al., 2024), and Gorilla (Patil et al., 2024) have demonstrated the ability to accurately use different tools like calculators and search engines. In parallel, strategies such as ToolkenGPT (Hao et al., 2023), METATOOL (Wang et al., 2024b), and IPR (Xiong et al., 2024) have been introduced to enhance the compatibility and effective use of various tools. Furthermore, agentic frameworks like WebMap (Spiegel & Horák, 2024), ReAct (Yao et al., 2023), Voyager (Wang et al., 2024a), Agent Reasoning (Wu et al., 2025), Middleware (Gu et al., 2024), and ViperGPT (Surís et al., 2023) enable models to perform complex, multi-step tasks. These applications range from online shopping and software repair to handling multimodal inputs and learning skills in interactive environments like Minecraft.
2.2 Tool-Use Benchmarks
@@
Numerous benchmarks have been developed to evaluate the tool-use capabilities of LLMs, each contributing different methodologies. Some efforts, such as ToolAlpaca (Tang et al., 2023) and NexusRaven (Srinivasan et al., 2023), have focused on automatic data generation to create large-scale evaluation sets. Another line of work has concentrated on creating diverse execution environments. For instance, Toolbench (Qin et al., 2024) provides a collection of non-executable APIs, while a more common strategy involves building sandboxed environments to simulate real-world interactions. BFCL-v3 (Patil et al., 2024), for example, implements codebases to simulate services like Twitter and mathematical calculators, and API-Bank (Li et al., 2023b) simulates common tasks such as sending emails or querying stock prices with mock data. Focusing on mobile-centric scenarios, ToolSandbox (Lu et al., 2025) simulates tools for managing device states like Wi-Fi, cellular service, and location service, while HammerBench (Wang et al., 2025) replicates API functionalities based on the features of 60 commercial applications. Other benchmarks have concentrated on specific domains, with ComplexFuncBench (Zhong et al., 2025) targeting travel-related tasks, $\tau$-bench (Yao et al., 2024) focusing on retail and airline scenarios, and ACEBench (Chen et al., 2025) covering four scenarios including food delivery and financial services. Fundamentally, these benchmarks assess model capabilities using simulated, simplified tools; their interaction logic and return values are based on fixed rules or templates, lacking the dynamism and uncertainty of real-world environments, and the complexity of the data they handle does not match that of practical applications.
@@
While previous work has attempted to connect LLMs with real-world APIs, such as in AutoFeedback (Liu et al., 2024; Song et al., 2023), these efforts often faced challenges due to complex and non-standardized interfaces. The Model Context Protocol (MCP) (Anthropic, 2024) was introduced to address this issue. MCP is a standardized protocol for LLM-tool interaction that enables dynamic discovery and orchestration of tools based on task requirements. As the protocol has matured, it has seen widespread adoption across the industry(OpenAI, 2025; cur, 2025; Anthropic, 2025). Consequently, recent work has begun to establish benchmarks on top of MCP. For instance, MCP-Zero (Fei et al., 2025) proposed a retrieval strategy for MCPs and collected an associated toolset, MCP-Tools. Meanwhile, MCPBench (Luo et al., 2025) focused on comparing the performance of different MCPs for web-search tasks, whereas MCP-RADAR (Gao et al., 2025) focused on mathematics and coding tasks and included a subset of general-purpose MCP scenarios.
3 MCPVerse Design
3.1 MCP Collection
@@
We collect a diverse set of MCPs from platforms such as Smithery, mcp.so, and ModelScope. These MCPs offer a wide range of functionalities, including local file operations, real-time flight booking, SQL querying and so on. As a Results, we carefully selected MCPs to be included in our benchmark based on the following criteria.
@@
In total, the selected $65$ MCPs expose more than $552$ individual tools. This number may fluctuate, as MCP providers have the discretion to add or deprecate tools over time.
3.2 Task Curation
@@
The MCP servers form a large and diverse action space, enabling the execution of a wide variety of task types. Based on this action space, we curate a benchmark consisting of $250$ tasks, categorized into two main types: Information Retrieval and System Operation, and further classified by their complexity, which is divided into three levels: L1, L2, and L3.
(a) Task Types
(b) Task Complexity Levels
(c) Task Time Sensitivity
Figure 2: Task distributions by (a) task type, (b) task complexity level, and (c) task time sensitivity.
Task Complexity Levels: L1 tasks are solvable by a single tool within $1$ or $2$ steps. L2 tasks require about $5$ steps and may involve a single tool or multiple tools working together. L3 tasks are relatively complex and demand multi-tool collaboration or in-depth application of a specific tool, typically requiring more than $5$ steps to complete.
Task Types: We categorize the tasks into two main types: Information Retrieval and System Operation. Information Retrieval tasks involve retrieving or querying specific types of information. These include subtypes such as: Geographical Information (location-based data), Financial Information (market trends, stock prices), Academic Research (retrieving academic papers or citations), and Hot News (real-time news updates). System Operation tasks include executing system commands (for example, cmd/shell), performing database operations, and handling files. Supported file types include Text Files (.txt), PDF Files (.pdf), Presentation Files (.ppt, .pptx), Document Files (.docx), and Spreadsheet Files (.xlsx).
@@
Curation Protocol: During the task construction phase, annotators (all undergraduate-level or above) select relevant MCPs from the MCP Hub according to the standards outlined earlier, then design tasks based on these MCPs. Annotators thoroughly review each tool definition within the MCPs and ensure the following: (1) A valid solution path exists within the selected set of MCPs. (2) The answers are objective and unambiguous to facilitate LLM-based scoring. (3) Tasks reflect real-world scenarios as closely as possible. (4) Tasks cannot be solved by the model without invoking external tools. Annotators use MCP client utilities (e.g., Cursor) to assist with question generation and answer retrieval. Each task is labeled with the following attributes: question, required MCPs, required tools, time sensitivity, complexity, task type, and ground truth. After initial construction, every task undergoes a comprehensive review. For time-sensitive tasks, custom scripts are employed to retrieve real-time answers for validation.
3.3 Evaluation Modes
Unlike many previous tool-use benchmarks that often incorporate a retrieval module or mount a target tool alongside a small set of unrelated tools (Fei et al., 2025; Li et al., 2023b; Qin et al., 2024), our objective is to significantly expand the toolset available to the LLM. This approach is motivated by two key reasons: (1) Recent advancements in large language models have significantly increased the context window, with models now capable of processing up to 1 million tokens. This alleviates the previous limitation where the toolset size was constrained by context length. (2) We aim to observe how LLMs perform in a large tool space, exploring new solution paths and exhibiting emergent behaviors or novel problem-solving strategies. A retriever trained on existing patterns might in- advertently hinder this crucial exploratory potential.
Based on this, we designed three distinct evaluation modes:
  • Oracle Mode: In this mode, we load only the minimal set of tools required to solve a given task. This minimal set is defined and annotated by the annotators, who, during the task construction process, verify the feasibility of using these tools to solve the task. As a result, the context length varies depending on the specific task and the tools required.